# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ 4.1 Future directions
Coupling of multiple scales and multi-body systems: Integrating multiple scales into nonlocal models allows researchers to bridge the gap between molecular-level interactions and tissue or organ-level behaviours. Future studies may focus on refining methods to couple various scales effectively, enabling more accurate predictions of biological processes at different hierarchical levels konieczny2023 . In the context of multi-scale porous media, such as biological tissues or cellular membranes, fractional diffusion models are used to describe how substances diffuse within the intricate network of interconnected pores. Traditional diffusion models are inadequate for accurately representing such complex structures, but fractional diffusion equations can effectively capture the anomalous transport observed in multiscale porous media sun2017 ; scheidweiler2020 . Furthermore, nonlocal interactions in multi-body systems are a fascinating research area, and their study has applications in diverse fields, including physics, materials science, and geophysics shaat2020 . As experimental techniques and computational methods continue to advance, our understanding of the role of nonlocal interactions in multi-body systems is expected to deepen, leading to new discoveries and applications in various disciplines.
@@
Cell migration: Nonlocal interactions in cell biology are a fascinating area of research, and they are essential in understanding how cells coordinate their activities in complex biological processes. Nonlocal models can be used to study the spatial and temporal aspects of immune responses, such as immune cell migration, pathogen spreading, and tissue repair. As experimental techniques and computational methods advance, the study of nonlocal interactions in cell biology will continue to provide valuable insights into the emergent behaviours and collective dynamics of cells in tissues and organisms buttenschon2020 . In addition, nonlocal models using fractional calculus could be used to capture the long-range communication and coordination among migrating cells, shedding light on the mechanisms governing collective cell behaviours.
@@
Hybrid models: Hybrid models can combine partial differential equations (PDEs) that describe spatial processes with ordinary differential equations (ODEs) that represent temporal dynamics. In this type of hybrid model, the local interactions represent the immediate influence of neighbouring elements, while nonlocal interactions account for long-range dependencies. These models could be used to study systems with spatial heterogeneity and temporal dynamics, such as cancer growth, morphogenesis, pattern formation, metastasis, competition, and predation in predator-prey interactions sherratt2016 ; volpert2009 ; bitsouni2018 . These models could be used further to model the spreading of signals, morphogens, or chemicals over extended distances in finding spatial patterns.
@@
Peridynamics: Peridynamics is a mathematical framework that extends classical continuum mechanics to describe the behaviour of materials with discontinuities and fractures. While peridynamics has been widely used in engineering and materials science, its applications in biology are still relatively new and evolving. It can be applied to model the mechanics of biological tissues, such as soft tissues (e.g., muscle, cartilage), hard tissues (e.g., bone), and invasion of cancer cells into surrounding tissues lejeune2021 . Furthermore, it can be employed to study the mechanics of biological membranes (e.g., lipid bilayers), cell migration, and the interactions between migrating cells and their surroundings.
@@
Epidemiological models: Nonlocal epidemiological models continue to be an important research area, especially in the context of emerging infectious diseases and the influence of spatial factors on disease transmission dynamics ruan2007 ; li2021 . Different types of physiological age structure characteristics of the population can be considered in the model to improve the predictions for infectious disease transmissions. As research in this field progresses, nonlocal models are expected to be increasingly important to guide public health interventions and inform disease control strategies in understanding infectious diseases, and computational capabilities also have to improve in parallel.
@@
Neural networks: In computational neuroscience, hybrid neural network models combine local, point-to-point synaptic connections with nonlocal, spatially extended interactions. This allows for the integration of both short-range and long-range interactions observed in real neural networks senk2022 . These models are employed to study brain function, learning, and memory processes. In addition, neurons in the brain have a multi-compartmental structure with dendrites, soma, and axons. Fractional-order differential equations can be used to model the electrical activity within neurons, accounting for the nonlocal interactions between different compartments magin2010 ; drapaca2016 . These models are essential for understanding how information is integrated and processed in complex neural networks.
@@
Fractional derivative models: Extending nonlocal models to the time domain using fractional calculus will enable modelling memory effects and history-dependent processes in biological systems. In the context of emerging infectious diseases, traditional local epidemiological models may fail to account for the impact of long-range dispersal and spatial heterogeneity. However, fractional epidemic models with nonlocal interactions can be used to study the spread of diseases in heterogeneous environments chen2021 . In addition, the folding and aggregation of proteins are complex processes influenced by long-range interactions between amino acids vascon2020 . In this case, the fractional models can be applied to study protein folding kinetics and aggregation dynamics to understand the underlying mechanisms of protein misfolding diseases, such as Alzheimer’s and Parkinson’s.
@@
Variational problems: There are many challenges and unique features associated with variational problems in understanding complex systems in biology, materials science, and social dynamics li2004 ; ramstead2019 . For instance, the analytical and numerical methods employed to tackle nonlocal variational problems, including the use of nonlocal differential operators and integral equations.
@@
Biosocial dynamics: The collective behaviour of animal groups, such as flocking birds or schooling fish, relies on nonlocal interactions that enable coordinated movements over large distances giardina2008 . Nonlocal models have been used to investigate these animal groups’ emergent patterns and self-organization, offering insights into their decision-making processes and survival strategies. Nevertheless, nonlocal models can offer a powerful framework for studying human social dynamics, allowing researchers to capture the complexity of interactions that extend beyond local connections, e.g., information spread, the epidemiological spread of behaviours, migration, long-term social evolution, etc.
Data-driven approaches: As biological systems are often complex and high-dimensional, future research might explore the development of efficient numerical methods and simulations tailored to handle nonlocal interactions. This includes optimization techniques for parameter estimation and sensitivity analysis to gain deeper insights into model behaviour. The integration of machine learning and artificial intelligence techniques with nonlocal models could enhance the predictive capabilities and understanding of complex biological phenomena. Future research may explore how artificial intelligence can assist in pattern recognition, model calibration, and model selection for nonlocal models. Developing efficient and scalable numerical methods for nonlocal models will be crucial to handling the computational demands of large-scale biological systems and high-dimensional data.
Brain: The brain is a highly complex organ with complicated networks and connections, and nonlocal effects can play a crucial role in understanding various aspects of brain function papo2024 . Neuronal networks often exhibit long-range connections, where the activity in one region can influence distant regions. Modelling with these nonlocal models helps to understand how information is processed and integrated across different brain regions. In addition, nonlocal diffusion processes can be used to model the spread of molecules, neurotransmitters, or electrical signals across the brain tissue.

### v2

@@ 4.1 Future directions
Coupling of multiple scales and multi-body systems: Integrating multiple scales into nonlocal models allows researchers to bridge the gap between molecular-level interactions and tissue or organ-level behaviours. Future studies may focus on refining methods to couple various scales effectively, enabling more accurate predictions of biological processes at different hierarchical levels [684]. In the context of multi-scale porous media, such as biological tissues or cellular membranes, fractional diffusion models are used to describe how substances diffuse within the intricate network of interconnected pores. Traditional diffusion models are inadequate for accurately representing such complex structures, but fractional diffusion equations can effectively capture the anomalous transport observed in multiscale porous media [685, 686]. Furthermore, nonlocal interactions in multi-body systems are a fascinating research area, and their study has applications in diverse fields, including physics, materials science, and geophysics [65]. As experimental techniques and computational methods continue to advance, our understanding of the role of nonlocal interactions in multi-body systems is expected to deepen, leading to discoveries and applications in various disciplines.
@@
Cell migration: Nonlocal interactions in cell biology are a fascinating area of research, and they are essential in understanding how cells coordinate their activities in complex biological processes. Nonlocal models can be used to study the spatial and temporal aspects of immune responses, such as immune cell migration, pathogen spreading, and tissue repair. As experimental techniques and computational methods advance, the study of nonlocal interactions in cell biology will continue to provide valuable insights into the emergent behaviours and collective dynamics of cells in tissues and organisms [687]. In addition, nonlocal models using fractional calculus could be used to capture the long-range communication and coordination among migrating cells, shedding light on the mechanisms governing collective cell behaviours.
@@
Hybrid models: Hybrid models can combine PDEs that describe spatial processes with ODEs that represent temporal dynamics. In this type of hybrid model, the local interactions represent the immediate influence of neighbouring elements, while nonlocal interactions account for long-range dependencies. These models could be used to study systems with spatial heterogeneity and temporal dynamics, such as cancer growth, morphogenesis, pattern formation, metastasis, competition, and predation in predator-prey interactions [54, 72, 379]. These models could be used further to model the spreading of signals, morphogens, or chemicals over extended distances in finding spatial patterns.
@@
Peridynamics: Peridynamics is a mathematical framework that extends classical continuum mechanics to describe the behaviour of materials with discontinuities and fractures. While peridynamics has been widely used in engineering and materials science, its applications in biology are still relatively new and evolving. It can be applied to model the mechanics of biological tissues, such as soft tissues (e.g., muscle, cartilage), hard tissues (e.g., bone), and invasion of cancer cells into surrounding tissues [68]. Furthermore, it can be employed to study the mechanics of biological membranes (e.g., lipid bilayers), cell migration, and the interactions between migrating cells and their surroundings.
@@
Epidemiological models: Nonlocal epidemiological models continue to be an important research area, especially in the context of emerging infectious diseases and the influence of spatial factors on disease transmission dynamics [134, 688]. Different types of physiological age structure characteristics of the population can be considered in the model to improve the predictions for infectious disease transmissions. As research in this field progresses, nonlocal models are expected to be increasingly important to guide public health interventions and inform disease control strategies in understanding infectious diseases, and computational capabilities also have to improve in parallel.
@@
Neural networks: In computational neuroscience, hybrid neural network models combine local, point-to-point synaptic connections with nonlocal, spatially extended interactions. This allows for the integration of both short-range and long-range interactions observed in real neural networks [689]. These models are employed to study brain function, learning, and memory processes. In addition, neurons in the brain have a multi-compartmental structure with dendrites, soma, and axons. Fractional-order differential equations can be used to model the electrical activity within neurons, accounting for the nonlocal interactions between different compartments [396, 690]. These models are essential for understanding how information is integrated and processed in complex neural networks.
@@
Fractional derivative models: Extending nonlocal models to the time domain using fractional calculus will enable modelling memory effects and history-dependent processes in biological systems. In the context of emerging infectious diseases, traditional local epidemiological models may fail to account for the impact of long-range dispersal and spatial heterogeneity. However, fractional epidemic models with nonlocal interactions can be used to study the spread of diseases in heterogeneous environments [44]. In addition, the folding and aggregation of proteins are complex processes influenced by long-range interactions between amino acids [691]. In this case, the fractional models can be applied to study protein folding kinetics and aggregation dynamics to understand the underlying mechanisms of protein misfolding diseases, such as Alzheimer’s and Parkinson’s.
@@
Variational problems: There are many challenges and unique features associated with variational problems in understanding complex systems in biology, materials science, and social dynamics [692, 693]. For instance, the analytical and numerical methods employed to tackle nonlocal variational problems, including the use of nonlocal differential operators and integral equations.
@@
Biosocial dynamics: The collective behaviour of animal groups, such as flocking birds or schooling fish, relies on nonlocal interactions that enable coordinated movements over large distances [694]. Nonlocal models have been used to investigate these animal groups’ emergent patterns and self-organization, offering insights into their decision-making processes and survival strategies. Nevertheless, nonlocal models can offer a powerful framework for studying human social dynamics, allowing researchers to capture the complexity of interactions that extend beyond local connections, e.g., information spread, the epidemiological spread of behaviours, migration, long-term social evolution, etc.
Data-driven approaches: As biological systems are often complex and high-dimensional, future research might explore the development of efficient numerical methods and simulations tailored to handle nonlocal interactions [695, 696]. This includes optimization techniques for parameter estimation and sensitivity analysis to gain deeper insights into model behaviour. The integration of machine learning and artificial intelligence techniques with nonlocal models could enhance the predictive capabilities and understanding of complex biological phenomena [314, 315, 316]. Future research may explore how artificial intelligence can assist in pattern recognition, model calibration, and model selection for nonlocal models. Developing efficient and scalable numerical methods for nonlocal models will be crucial to handling the computational demands of large-scale biological systems and high-dimensional data.
Brain: The brain is a highly complex organ with complicated networks and connections, and nonlocal effects can play a crucial role in understanding various aspects of brain function [619]. Neuronal networks often exhibit long-range connections, where the activity in one region can influence distant regions. Modelling with these nonlocal models helps to understand how information is processed and integrated across different brain regions. In addition, nonlocal diffusion processes can be used to model the spread of molecules, neurotransmitters, or electrical signals across the brain tissue.